OpenStudy (ttop0816):

please help!! For an acute angle x, tanxtan(x/2) cscx-1 secx-1 secx+1

hartnn (hartnn):

you know tan double angle formula ?

OpenStudy (ttop0816):

yes it is 2tan theta/1-tan^2theta!

hartnn (hartnn):

correct so tan x = tan 2(x/2) = ....

hartnn (hartnn):

tanx tan(x/2) = 2 tan^2 (x/2) / (1+ tan^2 (x/2)) see if you get this ?

hartnn (hartnn):

***** tanx tan(x/2) = 2 tan^2 (x/2) / (1- tan^2 (x/2))

hartnn (hartnn):

if you could get that, convert everything into sin and cos

OpenStudy (ttop0816):

oh okay wait let me convert it!

OpenStudy (ttop0816):

ummmm i got a really weird answer;; but i tried solving the whole thing and i got cscx-1

hartnn (hartnn):

is the question really tanx tan(x/2) ??

OpenStudy (ttop0816):

yes HAHAHAHA

hartnn (hartnn):

from the given choices, i think its sec x -1

OpenStudy (ttop0816):

oh okay! thank you~

hartnn (hartnn):

from the simple fact that tan x tan x/2 is 0 when x =0 and only sec x - 1 becomes 0 when x = 0